-- RunaBriguglio - 24 Nov 2010

new Intmat calibrated at ATT temp=10°
Tracknum modes iterations
20101124_131157 0-250 3
20101124_131321 251-500 3
20101124_131454 501-653 3

Flattening command:
tracknum electrical data optical data WFE
       
20101124_131157   154811 39
       
       

the flattened shape is 45nm RMS, maybe convection???? ATT tube temp is set -2° wrt mirror

noise meas:20101124_143000

new flattening with intmat 131157 but same result

new int mat, sampled again after the tubone has been set to 8° to avoid convection
Tracknum modes iterations notes

20101124_162523

20101124_170012

0-250 3 to be discarded, ff_pure_currents

20101124_162642 missing

20101124_164445

251-500 3  

20101124_162746

501-653 3  

the intmat is saved as 20101124_170012/flat_data654modes but the flattening is still bad

a new intmat is acquired but problems with the adam

power_off, power_on

Hp pos 246.3 1943.8 1953.8 45.26 99.3 0

new intmat, fixing some convection problems

20101124_191814 0-250 3  
20101124_191942 251-500 3  
20101124_192112 501-653 3  

data saved in 20101124_191814/flat_data654modes.sav

the flattening is still not converging. see data in folder flat10deg: meas1 and meas30sec_after are taken at 30 sec distance: the difference is 70nm RMS with large spots. the point is possibly that the system is flattening the spots sampled at t0 at t=t+60sec (after pseudoinverting the int mat) when they are no longer there.

tubone switched on at 203900

shell flattened with T0_fquir intmat (20101118_175313/20101122_211025_flat_data_654modes) WFE=38nm

electrical data=20101124_204914 optical data 20101124_204914_flat654 modes

then same step to verify if it is repeatable: new RMS is 47

applied int mat 191814 good result 30nm

electrical data saved in 20101124_210021/flattening_result, optical data saved in 210021_flat650modes

tested again but 44nm rms

test with nmeas=50 (=2sec, maybe too short....), rms=48, dara saved in 210800_50meas

test on the reapplication of the shape 20101124_210021/flattening_result

optical result is:20101124_211500_load_shape_20101124_210021 (force filtering is active, max force is 0.15)
Topic revision: r2 - 24 Nov 2010, RunaBriguglio
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback